OpenStudy (anonymous):

Use simple interest to find the ending balance. 1.) $600 at 14% for 3 years 2.) 50,300 at 15% for 6 years

OpenStudy (anonymous):

what's the formula for the simple interest?

OpenStudy (anonymous):

no response ??

OpenStudy (anonymous):

I = P r t

OpenStudy (anonymous):

I think so im sorry

OpenStudy (anonymous):

for the $600 how much interest it will accumulate after 3 years?

OpenStudy (anonymous):

600 * 14% * 3 = ?

OpenStudy (anonymous):

600 * 0.14 * 3 = 6 * 14 * 3 = $252 of interest earned after 3 years so the first question 's answer becomes balance = 600 + 252

OpenStudy (anonymous):

can u continue with the rest?